OpenStudy (anonymous):

Simplify: (x - 8)2 Simplify: (5x - 10)2 Simplify: (3a + 5b)2

OpenStudy (anonymous):

2x-16 10x-20 6a+10b

OpenStudy (anonymous):

\[(x-8)^{2} = x ^{2}-16x+64\] \[(5x-10)^{2}=25x ^{2}-100x+100\] \[(3a+5b)^{2}=9a ^{2}+30ab+25b ^{2}\]